OpenStudy (anonymous):

a golf driving range charges a flat fee of $20 to practice and then $5.75 for a bucket of balls. write an equation that models the charges (c) in terms of the number of buckets of balls (b) that you use.

jigglypuff314 (jigglypuff314):

Hello and Welcome to OpenStudy! :) for this we can use that slope intercept form y = mx + b where m = cost per bucket and b = flat rate do you know how to plug in from there? :)

OpenStudy (anonymous):

Flat rate of 20 dollars means that that price is constanct so 20 and 5,75 dollar per bucket of balls means it's a variable, it can change depending on how many buckets you want. So you can write the variable of 5,75 dollars as 5,75x. The x is a variable that is depending on how many buckets you get which changes your y. So as jigglypuff wrote down y=mx+b you can write it as c=20+5.75b